### v1

# Classifier-free graph diffusion for molecular property targeting

Abstract
This work focuses on the task of property targeting: that is, generating molecules conditioned on target chemical properties to expedite candidate screening for novel drug and materials development. DiGress is a recent diffusion model for molecular graphs whose distinctive feature is allowing property targeting through classifier-based (CB) guidance. While CB guidance may work to generate molecular-like graphs, we hint at the fact that its assumptions apply poorly to the chemical domain. Based on this insight we propose a classifier-free DiGress (FreeGress), which works by directly injecting the conditioning information into the training process. CF guidance is convenient given its less stringent assumptions and since it does not require to train an auxiliary property regressor, thus halving the number of trainable parameters in the model. We empirically show that our model yields up to 79% improvement in Mean Absolute Error with respect to DiGress on property targeting tasks on QM9 and ZINC-250k benchmarks. As an additional contribution, we propose a simple yet powerful approach to improve chemical validity of generated samples, based on the observation that certain chemical properties such as molecular weight correlate with the number of atoms in molecules.
1 Introduction
The problem of generating molecules with desired chemical properties is crucial to enable fast candidate screening in the early stages of drug (Dara et al. 2021) or novel materials (Liu et al. 2017) development. Although the advent of deep learning has achieved remarkable accomplishments in computational chemistry, this strategic endeavor remains an area of active research. In the past, much of the researchers’ focus has been on the task of property optimization, which is based on modifying an already generated molecule such that it acquires some solicited chemical property (Jin, Barzilay, and Jaakkola 2018). However, driven by the success of conditional generative models (Radford et al. 2021), an equally important – though slightly different – task has emerged recently: property targeting, that is, to natively generate molecules which satisfy pre-specified chemical desiderata (Sousa et al. 2021).
Among the plethora of deep generative models that have been developed in the last decade, some are based on textual representation of molecules such as SMILES (Weininger 1988) or SELFIES (Krenn et al. 2022), while others are based on generating the molecular graph directly. Regarding the latter category, which is of interest for this work, previous approaches have employed autoregressive generation conditioned on latent representations (e.g. Gómez-Bombarelli et al. 2018), as well as one-shot generation of the adjacency and node feature matrices of the molecular graph (e.g. De Cao and Kipf 2022).
Denoising Diffusion Probabilistic Models (DDPMs) (Ho, Jain, and Abbeel 2020) have been shown to achieve state-of-the-art performance in conditional generation, for example to synthesize images from user-provided textual guidance (Saharia et al. 2022). Due to their flexibility and excellent performance, DDPMs have been extended to the chemical domain and applied to with textual and graph-based approaches to tasks such as distribution learning, property optimization, and property targeting with promising results (Runcie and Mey 2023).
In particular, DiGress (Vignac et al. 2023) is one of the first successful applications of DDPMs for molecular generation. Under the hood, DiGress is based on a discrete diffusion process (Austin et al. 2021) which gradually applies noise to a molecular graph according to a transition matrix, while denoising is performed by a graph transformer network (Dwivedi and Bresson 2020). The most interesting feature of DiGress is the possibility to perform conditional generation for property targeting through classifier-based (CB) guidance (Dhariwal and Nichol 2021). Loosely speaking, CB guidance requires to train a separate classifier to predict the conditional information from noisy samples, and to inject the resulting gradients back into the reverse process to bias the generative process. While successful to some extent, CB guidance has been shown to be inherently limited by a) the necessity of training a separate property predictor, which defies the purpose of having a single conditional generative model in the first place, and b) the fact that the gradients of the property predictor are not always informative and may lead the generative process astray. Due to these limitations, in the current DDPMs practice classifier-free (CF) guidance is often preferred. The idea behind CF guidance is to directly incorporate the conditioning vector as input to train the conditional DDPM. With respect to CB guidance, CF guidance has been shown to enable more stable training and better generative performance in general (Ho and Salimans 2022).
With DiGress, and molecular generation specifically, CB guidance is further limited by the fact that the auxiliary model is a regressor whose predictions (the chemical properties) are assumed to be normally distributed even for noisy graphs which are chemically invalid. This motivates our first contribution, which consists of the development and implementation of CF guidance for DiGress, called FreeGress. Experimentally, we show that switching from CB to CF guidance is beneficial to improve at property targeting. In particular, we evaluated FreeGress against DiGress on the QM9 (Ramakrishnan et al. 2014) and ZINC-250k (Gómez-Bombarelli et al. 2018) datasets, where we queried the models to generate molecules with properties as close as possible to a target specification. Comparing the mean absolute error between the target properties and the properties of the generated molecules, FreeGress significantly outperformed DiGress, with improvements up to $79\%$ in the most favourable case. Besides improving performances, FreeGress does not require an auxiliary property regressor, which cuts the number of trainable parameters by half. Furthermore, guided by the observation that certain chemical properties relate to the molecular graph size (the number of atoms in the molecule), we also propose to improve the generative process by first learning the probability of sampling a certain number of nodes given the target property, and then using samples from this distribution to constrain the size of the graph to be generated. Through experiments, we show that this simple method allows to generate more chemically valid graphs without sacrificing performance.
2 Background and related works
2.1 Notation
For the purposes of this work, a molecular graph with $n$ atoms is a tuple $G=(\bm{X},\bm{\textsf{E}})$ where $\bm{X}\in\mathbb{R}^{n\times a}$ is a node matrix whose $i$-th row, indicated as $\bm{x}_{i}\in\mathbb{R}^{a}$, is the one-hot encoding of the atomic type of the $i$-th atom. Possible atomic types are specified by the set $\mathcal{X}$ with cardinality $a$. Similarly, $\bm{\textsf{E}}\in\mathbb{R}^{n\times n\times b}$ is an adjacency tensor (also called edge tensor) that jointly represents the molecular connectivity and the type of bond occurring between each pair of atoms. In practice, the $(i,j)$-th entry of E, indicated as $\bm{e}_{ij}\in\mathbb{R}^{b}$, contains the one-hot encoding of the bond type occurring between the $i$-th and $j$-th atoms. Possible bond types are specified by the set $\mathcal{E}$ with cardinality $b$, where the absence of a bond is itself considered a bond type. In the following we will use the term “node” interchangeably with “atom”, and “edge” interchangeably with “bond”. Conditioning vectors will be generally represented as vectors $\bm{y}\in\mathbb{R}^{d}$, where $d$ is usually a small integer (1 or 2 in our experiments).
2.2 Deep generative models for molecules
Historically, the first approach to use a deep generative model to produce unseen molecules has been proposed by Gómez-Bombarelli et al. (2018). Essentially, the model is a VAE where both the encoder and decoder are recurrent neural networks trained to reconstruct SMILES strings. In particular, the decoder is essentially an autoregressive SMILES language model capable of generating a molecule by sampling one SMILES character at a time. Although simplistic, this model still remains one of the most efficient. The model was designed to allow property optimization through Bayesian optimization in latent space. Over the years, several variants to address limitations such as low validity rates (Kusner, Paige, and Hernández-Lobato 2017) or the inability to handle more complex chemical substructures at the vocabulary level (Podda, Bacciu, and Micheli 2020) have been developed. Other methods related to this generative flavour use the alternative language of SELFIES instead, which is deemed to be more robust than SMILES as every possible character sequence defines a valid molecule. A different branch of methods is focused on generating the molecular graph directly. In this case, two possibilities have been explored: generating the graph autoregressively – either atom-by-atom (Li et al. 2018), edge-by-edge (You et al. 2018b; Goyal, Jain, and Ranu 2020; Bacciu and Podda 2021), or motif-by-motif (Jin, Barzilay, and Jaakkola 2020) – or directly predicting the adjacency and node feature matrices (De Cao and Kipf 2022). On graph-based models, property optimization has been performed with simple hill climbing in latent space (guided by a property predictor), up to iterative reinforcement learning approaches which assign higher rewards to chemically appealing molecules (You et al. 2018a). In particular, Jin, Barzilay, and Jaakkola (2020) propose to cast the property optimization problem as a molecular translation problem, where the difference between two molecules (one with suitable chemical properties, the other chemically similar but with poor chemical properties) is mapped to a Gaussian manifold. Latent vectors sampled from this prior can be used to modify unsuitable molecules into suitable ones. Lately, DDPMs that generate the molecular graph have started to be used for distribution learning as well as property optimization tasks. The primary application of DDPMs is 3D structure generation, which is useful to handle a different class of tasks such as ligand generation conditioned on the coordinates of a protein pocket (Baillif et al. 2023). Among DDPM-based models, DiGress was the first to be applied to 2D molecule generation, as well as one of the first to formalize the property targeting task.
2.3 DiGress: Denoising Diffusion for Graphs
Here we briefly describe DiGress (Vignac et al. 2023), a discrete DDPM for graphs which is the starting point of this study. Preliminarily, we provide a brief introduction to DDPM to set the notation and concepts for understanding our contribution.
Denoising Diffusion Probabilistic Models.
DDPMs consist of an untrained forward process $q(\bm{x}^{t}|\bm{x}^{t-1})$ paired with a parameterized reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$. The former iteratively perturbs the initial data point $\bm{x}^{0}$ to transform it into an $\bm{x}^{T}$ which is akin to Gaussian noise, while the second is trained to incrementally remove the noise from $\bm{x}^{T}$ until $\bm{x}^{0}$ is restored. Data generation from a trained DDPM is performed by sampling noise $\bm{\tilde{x}}^{T}\thicksim\mathcal{N}(0,\bm{I})$, where $\mathcal{N}$ is an isotropic Gaussian, and iteratively applying the denoising model for $T$ steps until a new datapoint $\bm{\tilde{x}}^{0}\thicksim q$ is obtained.
Conditioned generation in DDPMs is achieved by injecting a guidance vector, or guide $\bm{y}$, to obtain a conditioned reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$. CB guidance (Dhariwal and Nichol 2021) refactors the reverse process as $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})p_{\phi}(\bm{y}|\bm{x}^{t-1})$, where $p_{\phi}(\bm{y}|\bm{x}^{t-1})$ is an auxiliary classifier trained to predict the guide from a noisy version of the input. A scaled version of the gradient $\nabla_{\bm{x}^{t}}$ of $g_{\phi}$ is then added to the denoising gradient in order to magnify the conditioning signal. CB guidance is limited by the fact that it requires to train the auxiliary classifier, without the possibility of exploiting pre-trained models; moreover, since only few parts of the noisy $\bm{x}^{t}$ are actually useful to predict $\bm{y}$, the resulting gradient could yield undesirable directions in input space. In contrast, CF guidance (Ho and Salimans 2022) works by jointly optimizing $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$, the conditioned model, and $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$, the unconditioned model, at the same time. During sampling, the reverse process is computed as $(1+w)\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})-w\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$, which is a barycentric combination between the conditioned and unconditioned predictions with weight $w\in\mathbb{R}_{+}$.
DDPMs defined as above are not suited for discrete data such as graphs, since discrete distributions cannot be effectively corrupted using Gaussian noise. Instead, discrete DDPMs (Austin et al. 2021), including DiGress, inject noise in the data at time-step $t-1$ by multiplying it against a transition matrix $\bm{Q}^{t}$, where $[\bm{Q}^{t}]_{ij}=q(x^{t}=j|x^{t-1}=i)$, which loosely speaking specifies the probability of changing $\bm{x}$ from category $i$ to category $j$.
DiGress Forward process.
DiGress applies noise to a molecular graph $G$ at time-step $t$ according to the following process:
$q(G^{t}|G^{t-1})=(\bm{X}^{t-1}\bm{Q}^{t}_{\bm{X}},\bm{\textsf{E}}^{t-1}\bm{Q}^{t}_{\bm{\textsf{E}}}),$ (1)
where $\bm{Q}^{t}_{X}$ and $\bm{Q}^{t}_{\bm{\textsf{E}}}$ are transition matrices applied to the node matrix and the edge tensor, respectively. In particular, $\bm{Q}^{t}_{X}$ is chosen from the following distribution:
$\bm{Q}^{t}_{\bm{X}}=\alpha^{t}\bm{I}+\beta^{t}\bm{1}_{a}\bm{m}_{\bm{X}},$ (2)
where $\bm{m}_{\bm{X}}$ is a vector containing the marginal distribution of the node types in the training set, $\alpha^{t}$ is a noise scheduler, and $\beta^{t}=1-\alpha^{t}$. In practice, the transition probability to a certain node type is given by its frequency in the training set. The transition tensor $\bm{Q}^{t}_{\bm{\textsf{E}}}$ is chosen analogously. Unrolling the forward process equation, we can sample $G^{t}$ given $G^{0}$ in a single step as follows:
$q(G^{t}|G^{0})=(\bm{X}\overline{\bm{Q}}^{t}_{\bm{X}},\bm{\textsf{E}}\overline{\bm{Q}}^{t}_{\bm{\textsf{E}}}),$ (3)
where $\overline{\bm{Q}}^{t}_{\bm{X}}=\overline{\alpha}^{t}\bm{I}+\overline{\beta}^{t}\bm{1}_{a}\bm{m}_{\bm{X}}$, $\overline{\alpha}^{t}=\prod_{\tau=1}^{t}\alpha^{\tau}$, and $\overline{\beta}^{t}=1-\overline{\alpha}^{t}$. Notice that $\overline{\bm{Q}}^{t}_{\bm{\textsf{E}}}$ is defined analogously.
DiGress Reverse process.
DiGress reconstructs a molecular graph from a noisy sample with the following reverse process:
$p_{\theta}(G^{t-1}|G^{t}),$ (4)
which can be thought of as the joint distribution of the individual nodes and edges given the current graph. By assuming independence, we can rewrite the joint as a product of conditionals:
$\prod_{1\leq i\leq n}p_{\theta}(\bm{x}_{i}^{t-1}|G^{t})\prod_{1\leq i,j\leq n}p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t}).$ (5)
In turn, each node conditional is computed by marginalizing over the node types (resp. edge types for the edge conditionals). Specifically, we have:
$p_{\theta}(\bm{x}_{i}^{t-1}|G^{t})=\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}^{0}_{i}=\bm{x})\,f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t}),$ (6)
where $f_{\theta}$ is computed by a neural network trained to infer the node types of the true graph $G^{0}$ from a noisy version $G^{t}$. The edges marginal is computed similarly, by predicting the true edge types from their noisy intermediates. CB guidance for DiGress is achieved by specifying the conditioned reverse process as:
$p_{\theta}(G^{t-1}|G^{t})p_{\phi}(\bm{y}|G^{t-1}),$ (7)
where $p_{\phi}(\bm{y}|G^{t-1})\propto\exp(-\lambda\langle\triangledown_{G^{t}}||\bm{y}-g_{\phi}(G^{t})||^{2},G^{t-1}\rangle)$, $g_{\phi}$ is an auxiliary property regressor, and $\lambda$ is a hyper-parameter that magnifies the conditioning gradient.
Limitations.
While the development of unconditioned DiGress appears intuitive, the use of CB guidance for the conditional variant leads to design choices and assumptions that are unrealistic in the chemical domain. Firstly, CB guidance relies on an external predictor that learns chemical properties from noisy graphs, which implies that chemically invalid molecules are unnaturally related to valid ones by being assigned the same properties. Secondly, the distribution learned by the auxiliary property regressor is assumed to be Normal with mean $p(\bm{y}|G^{t})$, approximated by $g_{\phi}$ by minimizing a mean squared error objective. This assumption is unsupported by premise for most denoising steps (where $G^{t}$ usually represents an invalid molecule), since it implies a distribution of chemical properties for an object that does not even exist in chemical space. Moreover, even if two molecules with graphs $G^{t}$ and $G^{t-1}$ are both valid and similar, their chemical properties can be drastically different, since molecular property landscapes are in general non-smooth (Aldeghi et al. 2022), violating again the normality assumption. This discussion motivates our intention to develop CF guidance for DiGress with the objective of making the conditional process simpler and at the same time appropriate for the chemical context where it is applied.
3 Classifier-Free Graph Diffusion Process
In this section, we first work out how to incorporate CF guidance into the general architecture of DiGress. Then, we detail the specifics of the neural architecture implementing CF guidance. Finally, we describe a chemically-motivated variant of our approach based on predicting the number of nodes of the graph to be generated from the conditioning vector.
[…]
Figure 1: (a) The overall architecture of the proposed Neural Network. Notice that the guide $\bm{y}$ is now an input to the model. (b) The graph transformer layer for the input $\bm{X}$ (it is applied analogously to E and $\bm{u}$). (c) The self attention layer within the graph transformer layer. Notice that we superimposed ${}^{\prime}$ to the vectors indicate that they are network/layer outputs.
3.1 Classifier-free DiGress (FreeGress)
The forward noise process for our CF method can be straightforwardly inherited by discrete DDPM in general, and for the sake of this work from DiGress, as detailed in (1) and (2). The key impact of CF guidance in instead on the adaptive reverse process. To this end, we begin by rewriting the conditioned reverse process $p_{\theta}(G^{t-1}|G^{t},\bm{y})$ to incorporate $\bm{y}$ as follows:
$\prod_{1\leq i\leq n}p_{\theta}(\bm{x}_{i}^{t-1}|G^{t},\bm{y})\prod_{1\leq i,j\leq n}p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t},\bm{y}),$ (8)
which requires to reformulate the marginals from Eq. 6. Specifically, we can rewrite the nodes marginal along the lines of Gu et al. (2022) as:
$\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})\,f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t},\bm{y}),$ (9)
where $f_{\theta}$ is a neural network that predicts the true node types $\bm{x}^{0}$ using both the current noisy graph $G^{t}$ and the guide $\bm{y}$. Similarly, the edges marginal is expressed as:
$\sum_{\bm{e}\in\mathcal{E}}q(\bm{e}_{ij}^{t-1}|\bm{e}_{ij}^{t},\bm{e}_{ij}^{0}=\bm{e})\,f_{\theta}(\bm{e}_{ij}^{0}=\bm{e}|G^{t},\bm{y}).$ (10)
At this point, we have, in principle, a classifier-free DDPM. However, conditioning might not work as expected in this scenario, since the model could learn to completely ignore the guide $\bm{y}$. To prevent this undesirable behaviour, we leverage conditional dropout (Ho and Salimans 2022), where the conditioning vector $\bm{y}$ is replaced, with probability $p_{uncond}$, by a non-parametrized placeholder $\bm{\bar{y}}$. To summarize, the final vector $\bm{y}$ employed in the equations (and, consequently, in the neural network) is chosen as follows:
$\bm{y}\leftarrow\begin{cases}\bm{\bar{y}}&\text{with probability }p_{uncond}\\ \bm{y}&\text{with probability }1-p_{uncond}\end{cases}$ (11)
Intuitively, the placeholder signals the network to perform inference without guidance vector, as if the process was unconditioned. As a result, the model becomes capable of performing both conditioned and unconditioned sampling.
The major difference between (Ho and Salimans 2022) and our work is the fact that the former is designed for continuous signals, such as images, and their neural network is trained to predict the Gaussian noise $\epsilon$ corrupting the input signal $\bm{x}^{t}$ given both the latter and the conditioning vector $\bm{y}$. FreeGress, instead, is designed for discrete graphs and its neural network is trained to infer $G^{0}$ given the corrupted graph $\bm{G}^{t}$ and conditioning vector $\bm{y}$. To complete derivation of FreeGress, we build on Tang et al. (2023) and express $f_{\theta}$ as:
$f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t},\bm{y})=p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{y})\,p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t})^{s},$ (12)
where the first term is the standard reverse process, and the second forces the model to take the guide into account, with $s\in\mathbb{R}_{+}$ being its relative importance. This same objective can be expressed in log-probability space as:
$\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t})}+(s+1)\log p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t}).$ (13)
Lastly, using Bayes’ rule, the term $\log p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t})$ is decomposed as:
$\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{y})}-\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t})},$ (14)
which is computable since the model can output the unconditioned probability using the placeholder in place of the guide or, more formally, since ${p(\bm{x}^{0}_{i}|G^{t})}\approx{p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{\bar{y}}})$. The process is symmetric for the edges $\bm{e}\in\mathcal{E}$. Finally, to add more flexibility, we parametrize $\bm{\bar{y}}$ so that it is learned along with the parameters $\theta$.
Neural Architecture
In order to accommodate the explicit guidance, we modified the original DiGress architecture as shown in Figure 0(a). Basically, the network consists of a stack of Graph Transformer layers. In its original implementation, DiGress allows to augment the input matrices $\bm{X}$ and E with arbitrary node and edge features deemed useful (indicated with the $+$ superscript in Figure 0(a)). In addition, a vector $\bm{u}$ containing the current time step $t$ is added, which itself be enriched with graph-wise features. All these data are augmented with the guide $\bm{y}$ and processed by the stack of graph transformer layers. The graph transformer is displayed in Figure 0(b): at a high level, it resembles a standard transformer architecture, with a self attention module followed by dropout, residual connection, layer norm, and a feedforward stack. The self attention layer (shown in Figure 0(c)) first processes the (augmented) $\bm{X}$ matrix through the standard outer product and scaling typical of attention layers. Then, the result is used in conjunction with the edge tensor E as input to a FiLM layer (Perez et al. 2018) to produce unnormalized attention scores. The scores are passed through a softmax layer, combined once again with the values $\bm{X}$. The attention vectors are then flattened and mixed with the vector $\bm{u}$ through another FiLM layer, and then transformed linearly to obtain the prediction logits. The vector $\bm{u}$ is instead processed differently, as shown visually in the rightmost part of Figure 0(c). Specifically, $\bm{X}$ and E are first passed through independent PNA layers (Corso et al. 2020) and then summed together with a linear projection of $\bm{u}$ to form the final representation.
Loss function.
During training, FreeGress optimizes the following loss function for a given time step $t$:
$\delta\sum_{1\leq i\leq n}\mathrm{NLL}\left(\bm{x}_{i}^{0},f_{\theta}(\bm{x}_{i}^{t})\right)+\gamma\sum_{1\leq i,j\leq n}\mathrm{NLL}\left(\bm{e}_{ij}^{0},f_{\theta}(\bm{e}_{ij}^{t})\right),$
which is the negative log-likelihood (NLL) between the actual node (resp. edge) types and those predicted by the neural network described above. Parameters $\delta$ and $\gamma$ adjust the preference towards nodes or edges prediction.
Conditioned Inference
At each reverse process iteration, FreeGress generates a slightly denoised graph by first computing from the network $\log{p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{y})}$, the conditioned log probability, and $\log{p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{\bar{y}})}$, the unconditioned one. The final probability is obtained by exponentiating the sum of the two terms while giving an arbitrary (usually larger) weight $s$ to the conditioned probability:
$\log{p(\bm{x}^{0}_{i}|G^{t})}+(s+1)\,(\log{p(\bm{x}^{0}_{i}|G^{t},\bm{y})}-\log{p(\bm{x}^{0}_{i}|G^{t})}),$
where we omit exponentiation for conciseness, which corresponds to $f_{\theta}(\bm{x}^{0}=\bm{x}|G^{t},\bm{y})$ in Eq. 12. Then, the posteriors $q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})$ are computed $\forall\bm{x}\in\mathcal{X}$ and summed as in Eq. 9. This procedure is executed for each atom in the graph, and the resulting probabilities are sampled to obtain the nodes of the graph $G^{t-1}$. The process is symmetric for the graph edges.
3.2 Graph size inference
One limitation of diffusion models as a whole is the fact that the size (for graphs specifically, the number of nodes) of a sample cannot change during the denoising process. Some studies consider the possibility of inserting and removing elements from a generated, mono-dimensional sequence (Johnson et al. 2021) but, to the best of our knowledge, there are no similar works for graphs. In the original implementation of DiGress, the number of nodes the generated graph will have is sampled from the marginal distribution computed from the training and validation sets, respectively. While this is not an issue for unconditioned sampling, we observed that for conditioned sampling, a certain property (e.g. molecular weight) might be featured only by molecules with a specific number of atoms. To solve this problem, we propose that the number of nodes $n$ of the graph to be generated is sampled from $p_{\varepsilon}(n|\bm{y})$, which is parametrized as a neural network with two hidden layers and a softmax output layer. Here, the idea is to exploit the guide even before generation starts, by providing a graph size which is correlated with the requested property.
4 Experiments
Here, we detail the experimental analysis to evaluate FreeGress on property targeting tasks, that is, generating compounds which meet pre-specified molecular properties.
4.1 Dataset
Our experiments were conducted on QM9, a dataset of 133k small molecules (up to 9 heavy atoms) which was also part of the evaluation of DiGress, and ZINC-250k, a collection of 250k drug-like molecules selected from the ZINC dataset. For the latter, the molecules were first preprocessed by removing non-neutral charges and stereochemistry information, reducing the dataset size to 204k molecules.
4.2 Metrics and targets
All the experiments were performed using the following scheme. After we trained each model, we randomly sampled 100 molecules from the dataset and computed the desired target properties as vectors $\bm{y}$, which were used as conditioning vectors. Then, we used the vectors 10 times each to perform conditioned generation, for a total of 1000 generated molecules for each model. We computed the desired properties on the generated samples [^1] Using packages such as RDKit (Landrum 2016) and psi4 (Turney et al. 2012) and compared the results with the properties of the original samples. The metric chosen for the comparison is the Mean Absolute Error (MAE) between the target properties and the properties of the generated molecules:
$\frac{1}{1000}\sum_{i=1}^{100}\sum_{j=1}^{10}|\bm{y}_{i}-\widehat{\bm{y}}_{i,j}|,$
where $\bm{y}_{i}$ are the target properties of the $i$-th molecule from the dataset and $\widehat{\bm{y}}_{i,j}$ is the target property of the $j$-th molecule generated using the properties of the $i$-th molecule as guide. The targeted properties were dipole moment $\mu$ and the Highest Occupied Molecular Orbital (HOMO) for QM9; for ZINC-250k, we targeted log-partition coefficient (LogP) and Molecular Weight (MW).
Model Hyper-parameters Extra features HOMO $\downarrow$ $\bm{\mu}$downarrow HOMO + $\bm{\mu}$downarrow
DiGress Not provided No $0.56\pm 0.01$ $0.81\pm 0.04$ $0.87\pm 0.03$
FreeGress $p_{uncond}=0,s=5$ No $0.50\pm 0.05$ $0.99\pm 0.08$ $0.74\pm 0.06$
$p_{uncond}=0,s=5$ Yes $0.34\pm 0.03$ $0.95\pm 0.07$ $0.62\pm 0.04$
$p_{uncond}=0.1,s=5$ No $0.30\pm 0.04$ $0.60\pm 0.06$ $0.59\pm 0.07$
$p_{uncond}=0.1,s=5$ Yes $\mathbf{0.28\pm 0.04}$ $\mathbf{0.54\pm 0.06}$ $\mathbf{0.53\pm 0.07}$
$p_{uncond}=0.2,s=5$ No $0.31\pm 0.05$ $0.64\pm 0.07$ $0.62\pm 0.06$
$p_{uncond}=0.2,s=5$ Yes $0.36\pm 0.10$ $0.57\pm 0.07$ $0.55\pm 0.06$
Model Hyper-parameters LogP MW
MAE $\downarrow$ Valid $\uparrow$ MAE $\downarrow$ Valid $\uparrow$
DiGress $\lambda=50$ $1.15\pm 0.09$ 75% $73.40\pm 7.10$ 75%
$\lambda=500$ $0.60\pm 0.05$ 68% $60.30\pm 6.10$ 73%
$\lambda=1000$ $0.49\pm 0.05$ 65% $61.10\pm 5.90$ 69%
$\lambda=3000$ $0.52\pm 0.13$ 39% $64.10\pm 5.90$ 59%
FreeGress $p_{uncond}=0.1,s=2$ $\mathbf{0.15\pm 0.02}$ 85% $15.73\pm 2.74$ 58%
$p_{uncond}=0.1,s=2$, NI $0.16\pm 0.01$ 84% $\mathbf{15.18\pm 2.71}$ 75%
Table 1: Results on the QM9 (top) and ZINC-250k (bottom) datasets. Property targeting has been evaluated using Mean Absolute Error (see Section 4.2). HOMO: Highest Occupied Molecular Orbital; $\mathbf{\mu}$: dipole moment; $\lambda$: weight of the classier gradient for CB guidance; $s$: weight of conditioned inference for FreeGress; $p_{uncond}$: conditional dropout parameter for FreeGress; Valid: percentage of valid molecules among those sampled; NI: node inference (see Section 3.2).
4.3 Experimental details
QM9.
On QM9, we trained a set of FreeGress variants against the published results of DiGress. In the original article, DiGress was trained for the property targeting task without additional features, and hyper-parameters such as $\lambda$ (which regulates the impact of the regressor’s gradient) were not provided. FreeGress instances were trained with $s=5$, and with $p_{uncond}=\{0,0.1,0.2\}$ to study the effect of conditional dropout on the generative process. Moreover, we also trained different variants with and without additional features for $\bm{X}$ and $\bm{u}$, to understand if they provided any positive impact. Additional graph-wise features considered for this study were number of $k$-cycles with $k\in\{1,\ldots,6\}$, number of connected components, first 5 non-zero eigenvalues, molecular weight; among node-wise features, we considered the atom’s valency, the number of $k$-cycles a node belongs to, and an estimation of the biggest connected component size. FreeGress used similar hyper-parameters (5 graph transformer layers; embedding size of 256 for $\bm{X}$ and 128 for both E and $\bm{u}$; 8 attention heads; $T=500$, 1200 training epochs with batch size 512; Amsgrad optimizer (Reddi, Kale, and Kumar 2018) with learning rate of $2^{-4}$ and weight decay of $1^{-12}$; $\gamma=2\delta$) as DiGress for a fair comparison. Importantly, since FreeGress does not require an auxiliary regressor, all the variants are trained with half the number of parameters than DiGress (4.6 million vs. 9.2 million, approximately). All models required approximately seven hours to train on a nVidia A100 with 80GB of VRAM.
ZINC-250k.
For this dataset, we used a slightly different experimental setup since DiGress was not evaluated on this dataset previously. Specifically, we trained from scratch both DiGress and FreeGress with 12 layers and 1000 training epochs with batch size 256, while keeping the other hyper-parameters the same as for QM9. We used larger models since ZINC-250k includes bigger (in terms of number of atoms) and more diverse molecules. Training required approximately 100 hours for 1000 epochs each on the same machine as above. While FreeGress variants used approximately 16 million parameters, DiGress used approximately 32 million including the property regressor. We trained without additional features, since preliminary experiments have revealed that they had little to no influence on the performance, while at the same time largely increasing the training time. For DiGress, we evaluated different variants alternatives with $\lambda=\{50,500,1000,3000\}$. For FreeGress, we fixed $s=2$ and $p_{uncond}=0.1$ and evaluated one variant with the inference method proposed in Section 3.2, and one without. In the former case, $p_{\varepsilon}(n|\bm{y})$ was parameterized as a neural network with 2 layers with 512 units each, and ReLU activations.
Input $\mu$: 2.9326Estimated $\mu$: 2.8836
Input $\mu$: 4.1600Estimated $\mu$: 4.0071
Input HOMO: -6.7892Estimated HOMO: -6.6912
Input $\mu$: 2.9326Estimated $\mu$: 2.8836
Input $\mu$: 4.1600Estimated $\mu$: 4.0071
Input HOMO: -6.7892Estimated HOMO: -6.6912
Input HOMO: -7.4478Estimated HOMO: -7.3225
Figure 2: Curated molecules from the QM9 dataset.
5 Results
The experimental results are presented in Table 1. Novelty and uniqueness scores are not reported as they approached 100% in almost all cases. On the QM9 dataset (top table), we observe that FreeGress outperforms DiGress in almost every variant tested. Interestingly enough, even variants of FreeGress with $p_{uncond}=0$, i.e. those that use the guide without conditional dropout, achieve comparable performance with DiGress, with the only exception of the dipole moment $\mu$. Overall, the best performing variant of FreeGress ($p_{uncond}=0.1$, $s=5$, additional features) improves with respect of DiGress by 50% on the HOMO task, by 33% on the $\mu$ task, and by 39% on the combination of both. Similar improvements are achieved with slightly different combinations of the hyper-parameters, showing that CF guidance is robustly a better alternative to CB guidance for conditioned generation. Also, one should take into account that CF guidance does not require to train a separate auxiliary regressor, which saves up computational and memory resources. Four conditioned samples (two for each property targeting task) generated from the QM9 dataset are shown in Figure 2.
The analysis of performance in the ZINC-250k provides some additional interesting insights. Firstly, we observe a drop in validity of DiGress whenever too much weight is put on the auxiliary regressor by increasing $\lambda$. This is a clear signal that the gradient of the regressor is not providing useful directions in chemical space, which validates our belief that CB guidance, as formulated in DiGress, is based on unrealistic assumptions for the chemical domain. Secondly, we observe that FreeGress outperforms again DiGress on the two tasks by a great margin. Specifically, on the LogP task, FreeGress achieves a 69% improvement in MAE with respect to the best DiGress variant, while at the same time improving validity by 31%. On the MW task, the improvement in MAE is of about 79% with respect to the best DiGress variant. Thirdly, we observe that node inference has not a decisive impact in the LogP task. However, in the MW task, it improves the validity rate from 58% to 75% (on par with the best DiGress variant), while also slightly improving in terms of MAE. This aligns with our intuition: in fact, molecules with higher molecular weight are in general more likely to have more atoms than molecules with lower molecular weight. Thus, it makes sense, in this particular task, to having a mechanism to infer the number of nodes from the guide to bias the sampling process. Figure 2 shows 4 exemplay molecules (2 for each task) generated from the ZINC-250k dataset.
6 Conclusions
This work has developed and implemented FreeGress, a discrete DDPM for graphs with CF guidance, which allows to generate molecules complying to a pre-specified set of desired chemical properties. Through experiments, we have shown that CF guidance allows to generate better (more tailored to the specification) molecules than CB guidance without sacrificing chemical validity. Of course, there are still limitations to overcome, which we intend to address in future studies: for example, even if CF guidance allows to achieve good validity rates, the model is still far from achieving perfect validity. This might be addressed by designing the forward and reverse processes to use conditional transition matrices. Another way to increase the percentage of valid molecules generated without overfitting may be to train a model to work with molecular fragments instead of atoms, as it would likely reduce the amount of points in the molecular graph where an illegal structure is generated. We also have shown that inferring the number of nodes from the guide might be of help to generate graphs in some settings, especially when there is reason to believe that the guide is correlated with the graph’s size. As future work, we intend to validate this finding also in the more general setting of arbitrary graph generation.
Input LogP: 0.8472Estimated LogP: 0.8590
Input LogP: 2.1313Estimated LogP: 2.0127
Input MW: 401.13Estimated MW: 411.06
Input LogP: 0.8472Estimated LogP: 0.8590
Input LogP: 2.1313Estimated LogP: 2.0127
Input MW: 401.13Estimated MW: 411.06
Input MW: 195.02Estimated MW: 196.18
Figure 3: Curated molecules from the ZINC-250k dataset.

### v2

@@
-Classifier-free graph diffusion for molecular property targeting
+Matteo Ninniri,Marco Podda,Davide Bacciu
 Abstract
@@
 CF guidance is convenient given its less stringent assumptions and since it does not require to train an auxiliary property regressor, thus halving the number of trainable parameters in the model.
-We empirically show that our model yields up to 79% improvement in Mean Absolute Error with respect to DiGress on property targeting tasks on QM9 and ZINC-250k benchmarks.
-As an additional contribution, we propose a simple yet powerful approach to improve chemical validity of generated samples, based on the observation that certain chemical properties such as molecular weight correlate with the number of atoms in molecules.
+We empirically show that our model yields significant improvement in Mean Absolute Error with respect to DiGress on property targeting tasks on QM9 and ZINC-250k benchmarks.
+As an additional contribution, we propose a simple yet powerful approach to improve the chemical validity of generated samples, based on the observation that certain chemical properties such as molecular weight correlate with the number of atoms in molecules.
+Keywords:
+Graph Neural Networks Generative Models Deep Learning.
 1 Introduction
-The problem of generating molecules with desired chemical properties is crucial to enable fast candidate screening in the early stages of drug (Dara et al.
-2021) or novel materials (Liu et al.
-2017) development.
+00footnotetext: This preprint has not undergone peer review (when applicable) or any post-submission improvements or corrections.
+The Version of Record of this contribution is published "in Machine Learning and Knowledge Discovery in Databases.
+Research Track", and is available online at https://doi.org/10.1007/978-3-031-70359-1_19
+Generating molecules with desired chemical properties is crucial to enable fast candidate screening in the early stages of drug [4] or novel materials [14] development.
 Although the advent of deep learning has achieved remarkable accomplishments in computational chemistry, this strategic endeavor remains an area of active research.
-In the past, much of the researchers’ focus has been on the task of property optimization, which is based on modifying an already generated molecule such that it acquires some solicited chemical property (Jin, Barzilay, and Jaakkola 2018).
-However, driven by the success of conditional generative models (Radford et al.
-2021), an equally important – though slightly different – task has emerged recently: property targeting, that is, to natively generate molecules which satisfy pre-specified chemical desiderata (Sousa et al.
-2021).
-Among the plethora of deep generative models that have been developed in the last decade, some are based on textual representation of molecules such as SMILES (Weininger 1988) or SELFIES (Krenn et al.
-2022), while others are based on generating the molecular graph directly.
-Regarding the latter category, which is of interest for this work, previous approaches have employed autoregressive generation conditioned on latent representations (e.g.
-Gómez-Bombarelli et al.
-2018), as well as one-shot generation of the adjacency and node feature matrices of the molecular graph (e.g.
-De Cao and Kipf 2022).
-Denoising Diffusion Probabilistic Models (DDPMs) (Ho, Jain, and Abbeel 2020) have been shown to achieve state-of-the-art performance in conditional generation, for example to synthesize images from user-provided textual guidance (Saharia et al.
-2022).
-Due to their flexibility and excellent performance, DDPMs have been extended to the chemical domain and applied to with textual and graph-based approaches to tasks such as distribution learning, property optimization, and property targeting with promising results (Runcie and Mey 2023).
-In particular, DiGress (Vignac et al.
-2023) is one of the first successful applications of DDPMs for molecular generation.
-Under the hood, DiGress is based on a discrete diffusion process (Austin et al.
-2021) which gradually applies noise to a molecular graph according to a transition matrix, while denoising is performed by a graph transformer network (Dwivedi and Bresson 2020).
-The most interesting feature of DiGress is the possibility to perform conditional generation for property targeting through classifier-based (CB) guidance (Dhariwal and Nichol 2021).
+In the past, much of the researchers’ focus has been on the property optimization task, which is based on modifying a generated molecule such that it acquires some solicited chemical property [11].
+However, driven by the success of conditional generative models [16], an equally important – though slightly different – task has emerged recently: property targeting, that is, to natively generate molecules which satisfy pre-specified chemical desiderata [21].
+Denoising Diffusion Probabilistic Models (DDPMs) [9] have been shown to achieve state-of-the-art performance in conditional generation, for example to synthesize images from user-provided textual guidance [20].
+Due to their flexibility and excellent performance, DDPMs have been extended to the chemical domain, being applied to tasks such as distribution learning, property optimization, and property targeting with promising results [19].
+DiGress [24] is one of the first successful applications of DDPMs to molecular generation.
+Under the hood, DiGress is based on a discrete diffusion process [2] which gradually applies noise to a molecular graph according to a transition matrix, while denoising is performed by a graph transformer network [6].
+The most interesting feature of DiGress is the possibility to perform conditional generation for property targeting through classifier-based (CB) guidance [5].
 Loosely speaking, CB guidance requires to train a separate classifier to predict the conditional information from noisy samples, and to inject the resulting gradients back into the reverse process to bias the generative process.
-While successful to some extent, CB guidance has been shown to be inherently limited by a) the necessity of training a separate property predictor, which defies the purpose of having a single conditional generative model in the first place, and b) the fact that the gradients of the property predictor are not always informative and may lead the generative process astray.
-Due to these limitations, in the current DDPMs practice classifier-free (CF) guidance is often preferred.
+While successful to some extent, CB guidance has been shown to be inherently limited by i) the necessity of training a separate property predictor, which defies the purpose of having a single conditional generative model in the first place, and ii) the fact that the gradients of the property predictor are not always informative and may lead the generative process astray.
+Due to these limitations, classifier-free (CF) guidance for DDPMs is often preferred.
 The idea behind CF guidance is to directly incorporate the conditioning vector as input to train the conditional DDPM.
-With respect to CB guidance, CF guidance has been shown to enable more stable training and better generative performance in general (Ho and Salimans 2022).
-With DiGress, and molecular generation specifically, CB guidance is further limited by the fact that the auxiliary model is a regressor whose predictions (the chemical properties) are assumed to be normally distributed even for noisy graphs which are chemically invalid.
+With respect to CB guidance, CF guidance has been shown to enable more stable training and better generative performance in general [10].
+With DiGress for molecular generation specifically, CB guidance is further limited by the fact that the auxiliary model is a regressor whose predictions (the chemical properties) are assumed to be normally distributed even for noisy graphs which are chemically invalid.
 This motivates our first contribution, which consists of the development and implementation of CF guidance for DiGress, called FreeGress.
 Experimentally, we show that switching from CB to CF guidance is beneficial to improve at property targeting.
-In particular, we evaluated FreeGress against DiGress on the QM9 (Ramakrishnan et al.
-2014) and ZINC-250k (Gómez-Bombarelli et al.
-2018) datasets, where we queried the models to generate molecules with properties as close as possible to a target specification.
-Comparing the mean absolute error between the target properties and the properties of the generated molecules, FreeGress significantly outperformed DiGress, with improvements up to $79\%$ in the most favourable case.
-Besides improving performances, FreeGress does not require an auxiliary property regressor, which cuts the number of trainable parameters by half.
+In particular, we evaluated FreeGress against DiGress on the QM9 [17] and ZINC-250k [8] datasets, where we queried the models to generate molecules with properties as close as possible to a target specification.
+Comparing the mean absolute error between the target properties and the properties of the generated molecules, FreeGress significantly outperformed DiGress, with improvements up to $88\%$ in the most favourable case.
+Besides improving performance, FreeGress does not require an auxiliary property regressor, reducing the number of trainable parameters.
 Furthermore, guided by the observation that certain chemical properties relate to the molecular graph size (the number of atoms in the molecule), we also propose to improve the generative process by first learning the probability of sampling a certain number of nodes given the target property, and then using samples from this distribution to constrain the size of the graph to be generated.
 Through experiments, we show that this simple method allows to generate more chemically valid graphs without sacrificing performance.
+Our code and supplementary material are available at https://github.com/Asduffo/FreeGress.
 2 Background and related works
-2.1 Notation
-For the purposes of this work, a molecular graph with $n$ atoms is a tuple $G=(\bm{X},\bm{\textsf{E}})$ where $\bm{X}\in\mathbb{R}^{n\times a}$ is a node matrix whose $i$-th row, indicated as $\bm{x}_{i}\in\mathbb{R}^{a}$, is the one-hot encoding of the atomic type of the $i$-th atom.
+2.0.1 Notation.
+We represent a molecular graph with $n$ atoms as a tuple $G=(\bm{X},\bm{\textsf{E}})$ where $\bm{X}\in\mathbb{R}^{n\times a}$ is a node matrix whose $i$-th row, indicated as $\bm{x}_{i}\in\mathbb{R}^{a}$, is the one-hot encoding of the atomic type of the $i$-th atom.
 Possible atomic types are specified by the set $\mathcal{X}$ with cardinality $a$.
 Similarly, $\bm{\textsf{E}}\in\mathbb{R}^{n\times n\times b}$ is an adjacency tensor (also called edge tensor) that jointly represents the molecular connectivity and the type of bond occurring between each pair of atoms.
-In practice, the $(i,j)$-th entry of E, indicated as $\bm{e}_{ij}\in\mathbb{R}^{b}$, contains the one-hot encoding of the bond type occurring between the $i$-th and $j$-th atoms.
+The $(i,j)$-th entry of E, indicated as $\bm{e}_{ij}\in\mathbb{R}^{b}$, contains the one-hot encoding of the bond type occurring between the $i$-th and $j$-th atoms.
 Possible bond types are specified by the set $\mathcal{E}$ with cardinality $b$, where the absence of a bond is itself considered a bond type.
 In the following we will use the term “node” interchangeably with “atom”, and “edge” interchangeably with “bond”.
-Conditioning vectors will be generally represented as vectors $\bm{y}\in\mathbb{R}^{d}$, where $d$ is usually a small integer (1 or 2 in our experiments).
+Conditioning vectors will be generally represented as vectors $\bm{y}\in\mathbb{R}^{d}$, where $d$ is usually a small integer.
 2.1 Deep generative models for molecules
-Historically, the first approach to use a deep generative model to produce unseen molecules has been proposed by Gómez-Bombarelli et al.
-(2018).
+The first approach to use a deep generative model to produce unseen molecules has been proposed by [8].
 Essentially, the model is a VAE where both the encoder and decoder are recurrent neural networks trained to reconstruct SMILES strings.
-In particular, the decoder is essentially an autoregressive SMILES language model capable of generating a molecule by sampling one SMILES character at a time.
-Although simplistic, this model still remains one of the most efficient.
-The model was designed to allow property optimization through Bayesian optimization in latent space.
-Over the years, several variants to address limitations such as low validity rates (Kusner, Paige, and Hernández-Lobato 2017) or the inability to handle more complex chemical substructures at the vocabulary level (Podda, Bacciu, and Micheli 2020) have been developed.
 Other methods related to this generative flavour use the alternative language of SELFIES instead, which is deemed to be more robust than SMILES as every possible character sequence defines a valid molecule.
 A different branch of methods is focused on generating the molecular graph directly.
-In this case, two possibilities have been explored: generating the graph autoregressively – either atom-by-atom (Li et al.
-2018), edge-by-edge (You et al.
-2018b; Goyal, Jain, and Ranu 2020; Bacciu and Podda 2021), or motif-by-motif (Jin, Barzilay, and Jaakkola 2020) – or directly predicting the adjacency and node feature matrices (De Cao and Kipf 2022).
-On graph-based models, property optimization has been performed with simple hill climbing in latent space (guided by a property predictor), up to iterative reinforcement learning approaches which assign higher rewards to chemically appealing molecules (You et al.
-2018a).
-In particular, Jin, Barzilay, and Jaakkola (2020) propose to cast the property optimization problem as a molecular translation problem, where the difference between two molecules (one with suitable chemical properties, the other chemically similar but with poor chemical properties) is mapped to a Gaussian manifold.
-Latent vectors sampled from this prior can be used to modify unsuitable molecules into suitable ones.
+On graph-based models, property optimization has been performed with simple hill climbing in latent space (guided by a property predictor), up to iterative reinforcement learning approaches which assign higher rewards to chemically appealing molecules [25].
 Lately, DDPMs that generate the molecular graph have started to be used for distribution learning as well as property optimization tasks.
-The primary application of DDPMs is 3D structure generation, which is useful to handle a different class of tasks such as ligand generation conditioned on the coordinates of a protein pocket (Baillif et al.
-2023).
-Among DDPM-based models, DiGress was the first to be applied to 2D molecule generation, as well as one of the first to formalize the property targeting task.
+2.2 Denoising Diffusion Probabilistic Models
+DDPMs consist of an untrained forward process $q(\bm{x}^{t}|\bm{x}^{t-1})$ and a parameterized reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$.
+The former iteratively perturbs the initial data point $\bm{x}^{0}$ to transform it into an $\bm{x}^{T}$ akin to Gaussian noise, while the second is trained to incrementally remove the noise from $\bm{x}^{T}$ until $\bm{x}^{0}$ is restored.
+Generation from a trained DDPM amounts to sample $\bm{\tilde{x}}^{T}\thicksim\mathcal{N}(0,\bm{I})$ from an isotropic Gaussian $\mathcal{N}$ and iteratively applying the denoising model for $T$ steps until a new sample $\bm{\tilde{x}}^{0}\thicksim q$ is obtained.
+Conditioned generation in DDPMs is achieved by injecting a guidance vector, or guide $\bm{y}$, to obtain a conditioned reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$.
+DDPMs defined as above are not suited for graphs, since Gaussian noising does not work effectively on discrete distributions.
+2.2.1 Classifier-based guidance
+CB guidance [5] refactors the reverse process as $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})p_{\phi}(\bm{y}|\bm{x}^{t-1})$, where $p_{\phi}(\bm{y}|\bm{x}^{t-1})$ employs an auxiliary classifier trained to predict the guide from a noisy version of the input.
+A scaled version of the regressor gradient $\nabla_{\bm{x}^{t}}$ is used to magnify the conditioning signal.
+CB guidance is limited by the fact that it requires to train the auxiliary classifier, without the possibility of exploiting pre-trained models; moreover, since only few parts of the noisy $\bm{x}^{t}$ are actually useful to predict $\bm{y}$, the resulting gradient could yield undesirable directions in input space.
+2.2.2 Classifier-free guidance
+Differently from the CB approach, CF guidance [10] jointly optimizes $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$, the conditioned model, and $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$, the unconditioned model, at the same time.
+During sampling, the reverse process is computed as the barycentric combination $(1+w)\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})-w\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$ between the conditioned and unconditioned predictions with weight $w\in\mathbb{R}_{+}$.
 2.3 DiGress: Denoising Diffusion for Graphs
-Here we briefly describe DiGress (Vignac et al.
-2023), a discrete DDPM for graphs which is the starting point of this study.
-Preliminarily, we provide a brief introduction to DDPM to set the notation and concepts for understanding our contribution.
-Denoising Diffusion Probabilistic Models.
-DDPMs consist of an untrained forward process $q(\bm{x}^{t}|\bm{x}^{t-1})$ paired with a parameterized reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$.
-The former iteratively perturbs the initial data point $\bm{x}^{0}$ to transform it into an $\bm{x}^{T}$ which is akin to Gaussian noise, while the second is trained to incrementally remove the noise from $\bm{x}^{T}$ until $\bm{x}^{0}$ is restored.
-Data generation from a trained DDPM is performed by sampling noise $\bm{\tilde{x}}^{T}\thicksim\mathcal{N}(0,\bm{I})$, where $\mathcal{N}$ is an isotropic Gaussian, and iteratively applying the denoising model for $T$ steps until a new datapoint $\bm{\tilde{x}}^{0}\thicksim q$ is obtained.
-Conditioned generation in DDPMs is achieved by injecting a guidance vector, or guide $\bm{y}$, to obtain a conditioned reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$.
-CB guidance (Dhariwal and Nichol 2021) refactors the reverse process as $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})p_{\phi}(\bm{y}|\bm{x}^{t-1})$, where $p_{\phi}(\bm{y}|\bm{x}^{t-1})$ is an auxiliary classifier trained to predict the guide from a noisy version of the input.
-A scaled version of the gradient $\nabla_{\bm{x}^{t}}$ of $g_{\phi}$ is then added to the denoising gradient in order to magnify the conditioning signal.
-CB guidance is limited by the fact that it requires to train the auxiliary classifier, without the possibility of exploiting pre-trained models; moreover, since only few parts of the noisy $\bm{x}^{t}$ are actually useful to predict $\bm{y}$, the resulting gradient could yield undesirable directions in input space.
-In contrast, CF guidance (Ho and Salimans 2022) works by jointly optimizing $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$, the conditioned model, and $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$, the unconditioned model, at the same time.
-During sampling, the reverse process is computed as $(1+w)\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})-w\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$, which is a barycentric combination between the conditioned and unconditioned predictions with weight $w\in\mathbb{R}_{+}$.
-DDPMs defined as above are not suited for discrete data such as graphs, since discrete distributions cannot be effectively corrupted using Gaussian noise.
-Instead, discrete DDPMs (Austin et al.
-2021), including DiGress, inject noise in the data at time-step $t-1$ by multiplying it against a transition matrix $\bm{Q}^{t}$, where $[\bm{Q}^{t}]_{ij}=q(x^{t}=j|x^{t-1}=i)$, which loosely speaking specifies the probability of changing $\bm{x}$ from category $i$ to category $j$.
-DiGress Forward process.
-DiGress applies noise to a molecular graph $G$ at time-step $t$ according to the following process:
-$q(G^{t}|G^{t-1})=(\bm{X}^{t-1}\bm{Q}^{t}_{\bm{X}},\bm{\textsf{E}}^{t-1}\bm{Q}^{t}_{\bm{\textsf{E}}}),$ (1)
-where $\bm{Q}^{t}_{X}$ and $\bm{Q}^{t}_{\bm{\textsf{E}}}$ are transition matrices applied to the node matrix and the edge tensor, respectively.
-In particular, $\bm{Q}^{t}_{X}$ is chosen from the following distribution:
+Here we briefly describe DiGress [24], a discrete DDPM for graphs which is the starting point of this study.
+In the forward process, DiGress injects noise in the data at time-step $t-1$ by multiplying it with a transition matrix, which loosely speaking specifies the probability of transitioning from one node type to another.
+The forward process is applied until the training graph is completely corrupted.
+The reverse process implemented by DiGress is specified as $p_{\theta}(G^{t-1}|G^{t})$, which is factored (by assuming node and edge independence) as a product of conditionals of the individual nodes and edges given the current graph.
+In turn, each node conditional is obtained by marginalizing over the node types (resp.
+edge types for the edge conditionals).
+Specifically, the nodes (resp.
+edge) marginal is computed by predicting the true node (resp.
+edge) types from their noisy intermediates.
+For conditioned generation, DiGress employs CB guidance.
+Specifically, the conditioned reverse process is formulated as:
+$p_{\theta}(G^{t-1}|G^{t})p_{\phi}(\bm{y}|G^{t-1}),$ (1)
+where $p_{\phi}(\bm{y}|G^{t-1})\propto\exp(-\lambda\langle\triangledown_{G^{t}}||\bm{y}-g_{\phi}(G^{t})||^{2},G^{t-1}\rangle)$, $g_{\phi}$ is a property regressor, and $\lambda$ is a hyper-parameter that magnifies the conditional gradient.
+2.3.1 Limitations
+While the development of unconditioned DiGress appears intuitive, the use of CB guidance for the conditional variant leads to design choices and assumptions that are unrealistic in the chemical domain.
+Firstly, CB guidance relies on an external predictor that learns chemical properties from noisy graphs, which implies that chemically invalid molecules are unnaturally related to valid ones by being assigned the same properties.
+Secondly, the distribution learned by the auxiliary property regressor is assumed to be Normal with mean $p(\bm{y}|G^{t})$, approximated by $g_{\phi}$ by minimizing a mean squared error objective.
+This assumption is unsupported by premise for most denoising steps (where $G^{t}$ usually represents an invalid molecule), since it implies a distribution of chemical properties for an object that does not even exist in chemical space.
+Moreover, even if two molecules with graphs $G^{t}$ and $G^{t-1}$ are both valid and similar, their chemical properties can be drastically different, since molecular property landscapes are in general non-smooth [1], violating again the normality assumption.
+This discussion motivates our intention to develop CF guidance for DiGress with the objective of making the conditional process simpler and at the same time appropriate for the chemical context where it is applied.
+[…]
+Figure 1: A depiction of FreeGress.
+The forward process, which gradually corrupts a molecule into a random graph, goes from left to the right.
+The reverse process, which denoises the original graph, goes from right to left.
+Note that the reverse process allows for a conditioning vector $\bm{y}$ and a number of nodes $n$ sampled from a trained neural network $p_{\varepsilon}$.
+3 Classifier-Free Graph Diffusion
+This section presents our first contribution, a classifier-free DDPM for conditioned molecular generation named FreeGress. A high-level summarization of the proposed model is sketched in Figure 1.
+3.0.1 Forward process
+The forward process of FreeGress is analogous to that of DiGress, and essentially consists of applying $T$ noising steps to an input molecular graph $G$ according to the following process:
+$q(G^{t}|G^{t-1})=(\bm{X}^{t-1}\bm{Q}^{t}_{\bm{X}},\bm{\textsf{E}}^{t-1}\bm{Q}^{t}_{\bm{\textsf{E}}}),\;1\leq t\leq T$ (2)
+where $\bm{Q}^{t}_{X}$ and $\bm{Q}^{t}_{\bm{\textsf{E}}}$ are transition matrices applied to the node matrix and the edge tensor, respectively. In particular, $\bm{Q}^{t}_{X}$ is formulated as:
 $\bm{Q}^{t}_{\bm{X}}=\alpha^{t}\bm{I}+\beta^{t}\bm{1}_{a}\bm{m}_{\bm{X}},$ (3)
@@
 Notice that $\overline{\bm{Q}}^{t}_{\bm{\textsf{E}}}$ is defined analogously.
-DiGress Reverse process.
-DiGress reconstructs a molecular graph from a noisy sample with the following reverse process:
-$p_{\theta}(G^{t-1}|G^{t}),$ (4)
-which can be thought of as the joint distribution of the individual nodes and edges given the current graph. By assuming independence, we can rewrite the joint as a product of conditionals:
-$\prod_{1\leq i\leq n}p_{\theta}(\bm{x}_{i}^{t-1}|G^{t})\prod_{1\leq i,j\leq n}p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t}).$ (5)
-In turn, each node conditional is computed by marginalizing over the node types (resp. edge types for the edge conditionals). Specifically, we have:
-$p_{\theta}(\bm{x}_{i}^{t-1}|G^{t})=\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}^{0}_{i}=\bm{x})\,f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t}),$ (6)
-where $f_{\theta}$ is computed by a neural network trained to infer the node types of the true graph $G^{0}$ from a noisy version $G^{t}$.
-The edges marginal is computed similarly, by predicting the true edge types from their noisy intermediates.
-CB guidance for DiGress is achieved by specifying the conditioned reverse process as:
-$p_{\theta}(G^{t-1}|G^{t})p_{\phi}(\bm{y}|G^{t-1}),$ (7)
-where $p_{\phi}(\bm{y}|G^{t-1})\propto\exp(-\lambda\langle\triangledown_{G^{t}}||\bm{y}-g_{\phi}(G^{t})||^{2},G^{t-1}\rangle)$, $g_{\phi}$ is an auxiliary property regressor, and $\lambda$ is a hyper-parameter that magnifies the conditioning gradient.
-Limitations.
-While the development of unconditioned DiGress appears intuitive, the use of CB guidance for the conditional variant leads to design choices and assumptions that are unrealistic in the chemical domain.
-Firstly, CB guidance relies on an external predictor that learns chemical properties from noisy graphs, which implies that chemically invalid molecules are unnaturally related to valid ones by being assigned the same properties.
-Secondly, the distribution learned by the auxiliary property regressor is assumed to be Normal with mean $p(\bm{y}|G^{t})$, approximated by $g_{\phi}$ by minimizing a mean squared error objective.
-This assumption is unsupported by premise for most denoising steps (where $G^{t}$ usually represents an invalid molecule), since it implies a distribution of chemical properties for an object that does not even exist in chemical space.
-Moreover, even if two molecules with graphs $G^{t}$ and $G^{t-1}$ are both valid and similar, their chemical properties can be drastically different, since molecular property landscapes are in general non-smooth (Aldeghi et al.
-2022), violating again the normality assumption.
-This discussion motivates our intention to develop CF guidance for DiGress with the objective of making the conditional process simpler and at the same time appropriate for the chemical context where it is applied.
-3 Classifier-Free Graph Diffusion Process
-In this section, we first work out how to incorporate CF guidance into the general architecture of DiGress.
-Then, we detail the specifics of the neural architecture implementing CF guidance.
-Finally, we describe a chemically-motivated variant of our approach based on predicting the number of nodes of the graph to be generated from the conditioning vector.
+3.0.2 Reverse Process
+The key impact of CF guidance is instead on the adaptive reverse process. To this end, we begin by rewriting the conditioned reverse process as:
+$p_{\theta}(G^{t-1}|G^{t},\bm{y}),$ (5)
+such that it now incorporates the guide $\bm{y}$ as input. Then, we assume independence between nodes and edges and factorize as the following product of conditionals:
+$p_{\theta}(G^{t-1}|G^{t},\bm{y})=\prod_{1\leq i\leq n}p_{\theta}(\bm{x}_{i}^{t-1}|G^{t},\bm{y})\prod_{1\leq i,j\leq n}p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t},\bm{y}).$ (6)
+Following [7], we rewrite the nodes marginal as:
+$p_{\theta}(\bm{x}_{i}^{t-1}|G^{t},\bm{y})=\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})\,f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t},\bm{y}),$ (7)
+where $f_{\theta}$ is a neural network that predicts the true node types $\bm{x}^{0}$ using both the current noisy graph $G^{t}$ and the guide $\bm{y}$. Similarly, the edges marginal is:
+$p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t},\bm{y})=\sum_{\bm{e}\in\mathcal{E}}q(\bm{e}_{ij}^{t-1}|\bm{e}_{ij}^{t},\bm{e}_{ij}^{0}=\bm{e})\,f_{\theta}(\bm{e}_{ij}^{0}=\bm{e}|G^{t},\bm{y}).$ (8)
+At this point, we have, in principle, a classifier-free DDPM for graphs.
+However, conditioning might not work as expected in this scenario, since the model could learn to completely ignore the guide $\bm{y}$.
+This behaviour has been observed, e.g., by [22].
+To prevent this, we leverage conditional dropout [10], where the conditioning vector $\bm{y}$ is replaced, with probability $\rho$, by a non-parametrized placeholder $\bm{\bar{y}}$.
+To summarize, the final vector $\bm{y}$ employed in the equations (and, consequently, in the neural network) is chosen as follows:
+$\bm{y}\leftarrow\begin{cases}\bm{\bar{y}}&\text{with probability }\rho\\ \bm{y}&\text{with probability }1-\rho\end{cases}$ (9)
+Intuitively, the placeholder signals the network to perform inference without a guidance vector, as if the process was unconditioned.
+As a result, the model becomes capable of performing both conditioned and unconditioned sampling.
+To complete the derivation of FreeGress, we build on [10] and express $f_{\theta}$ as:
+$p_{\theta}(\bm{x}^{0}_{i}|G^{t})+s(p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{y})-p_{\theta}(\bm{x}^{0}_{i}|G^{t}))$ (10)
+which is computable since the model can output the unconditioned probability using the placeholder in place of the guide or, more formally, since ${p(\bm{x}^{0}_{i}|G^{t})}\approx{p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{\bar{y}}})$.
+The process is symmetric for the edges $\bm{e}\in\mathcal{E}$.
+Finally, to add more flexibility, we follow [22] and we parametrize $\bm{\bar{y}}$ so that it is trainable.
+The major difference between [10] and our work is the fact that the former is designed for continuous signals, such as images, and their neural network is trained to predict the Gaussian noise $\epsilon$ corrupting the input signal $\bm{x}^{t}$ given both the latter and the conditioning vector $\bm{y}$.
+FreeGress, instead, is designed for discrete graphs and its neural network is trained to infer $G^{0}$ given the corrupted graph $G^{t}$ and conditioning vector $\bm{y}$.
+An alternative formula, proposed by [22], is shown in Equation 11:
+$\log{p(\bm{x}^{0}_{i}|G^{t})}+s(\log{p(\bm{x}^{0}_{i}|G^{t},\bm{y})}-\log{p(\bm{x}^{0}_{i}|G^{t})}).$ (11)
+Note how setting the hyper-parameter $s$ to zero is equal to disabling the conditional dropout altogether.
 […]
@@
 […]
-Figure 1: (a) The overall architecture of the proposed Neural Network.
-Notice that the guide $\bm{y}$ is now an input to the model.
-(b) The graph transformer layer for the input $\bm{X}$ (it is applied analogously to E and $\bm{u}$).
+Figure 2: (a) The overall architecture of FreeGress.
+Notice that the guide $\bm{y}$ is an input to the model.
+(b) The graph transformer layer.
+$\bm{V}$ represents the triple $(\bm{X},\bm{\textsf{E}},\bm{u})$.
 (c) The self attention layer within the graph transformer layer.
-Notice that we superimposed ${}^{\prime}$ to the vectors indicate that they are network/layer outputs.
-3.1 Classifier-free DiGress (FreeGress)
-The forward noise process for our CF method can be straightforwardly inherited by discrete DDPM in general, and for the sake of this work from DiGress, as detailed in (1) and (2).
-The key impact of CF guidance in instead on the adaptive reverse process.
-To this end, we begin by rewriting the conditioned reverse process $p_{\theta}(G^{t-1}|G^{t},\bm{y})$ to incorporate $\bm{y}$ as follows:
-$\prod_{1\leq i\leq n}p_{\theta}(\bm{x}_{i}^{t-1}|G^{t},\bm{y})\prod_{1\leq i,j\leq n}p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t},\bm{y}),$ (8)
-which requires to reformulate the marginals from Eq. 6. Specifically, we can rewrite the nodes marginal along the lines of Gu et al. (2022) as:
-$\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})\,f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t},\bm{y}),$ (9)
-where $f_{\theta}$ is a neural network that predicts the true node types $\bm{x}^{0}$ using both the current noisy graph $G^{t}$ and the guide $\bm{y}$. Similarly, the edges marginal is expressed as:
-$\sum_{\bm{e}\in\mathcal{E}}q(\bm{e}_{ij}^{t-1}|\bm{e}_{ij}^{t},\bm{e}_{ij}^{0}=\bm{e})\,f_{\theta}(\bm{e}_{ij}^{0}=\bm{e}|G^{t},\bm{y}).$ (10)
-At this point, we have, in principle, a classifier-free DDPM.
-However, conditioning might not work as expected in this scenario, since the model could learn to completely ignore the guide $\bm{y}$.
-To prevent this undesirable behaviour, we leverage conditional dropout (Ho and Salimans 2022), where the conditioning vector $\bm{y}$ is replaced, with probability $p_{uncond}$, by a non-parametrized placeholder $\bm{\bar{y}}$.
-To summarize, the final vector $\bm{y}$ employed in the equations (and, consequently, in the neural network) is chosen as follows:
-$\bm{y}\leftarrow\begin{cases}\bm{\bar{y}}&\text{with probability }p_{uncond}\\ \bm{y}&\text{with probability }1-p_{uncond}\end{cases}$ (11)
-Intuitively, the placeholder signals the network to perform inference without guidance vector, as if the process was unconditioned.
-As a result, the model becomes capable of performing both conditioned and unconditioned sampling.
-The major difference between (Ho and Salimans 2022) and our work is the fact that the former is designed for continuous signals, such as images, and their neural network is trained to predict the Gaussian noise $\epsilon$ corrupting the input signal $\bm{x}^{t}$ given both the latter and the conditioning vector $\bm{y}$.
-FreeGress, instead, is designed for discrete graphs and its neural network is trained to infer $G^{0}$ given the corrupted graph $\bm{G}^{t}$ and conditioning vector $\bm{y}$.
-To complete derivation of FreeGress, we build on Tang et al.
-(2023) and express $f_{\theta}$ as:
-$f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t},\bm{y})=p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{y})\,p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t})^{s},$ (12)
-where the first term is the standard reverse process, and the second forces the model to take the guide into account, with $s\in\mathbb{R}_{+}$ being its relative importance.
-This same objective can be expressed in log-probability space as:
-$\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t})}+(s+1)\log p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t}).$ (13)
-Lastly, using Bayes’ rule, the term $\log p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t})$ is decomposed as:
-$\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{y})}-\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t})},$ (14)
-which is computable since the model can output the unconditioned probability using the placeholder in place of the guide or, more formally, since ${p(\bm{x}^{0}_{i}|G^{t})}\approx{p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{\bar{y}}})$.
-The process is symmetric for the edges $\bm{e}\in\mathcal{E}$.
-Finally, to add more flexibility, we parametrize $\bm{\bar{y}}$ so that it is learned along with the parameters $\theta$.
-Neural Architecture
-In order to accommodate the explicit guidance, we modified the original DiGress architecture as shown in Figure 0(a).
-Basically, the network consists of a stack of Graph Transformer layers.
-In its original implementation, DiGress allows to augment the input matrices $\bm{X}$ and E with arbitrary node and edge features deemed useful (indicated with the $+$ superscript in Figure 0(a)).
-In addition, a vector $\bm{u}$ containing the current time step $t$ is added, which itself be enriched with graph-wise features.
-All these data are augmented with the guide $\bm{y}$ and processed by the stack of graph transformer layers.
+Notice that we superimposed to the vectors indicate that they are network/layer outputs.
+3.0.3 Neural Architecture.
+Figure 2(a) depicts the neural architecture of our approach integrating explicit guidance.
+The network consists of a stack of Graph Transformer layers.
+The input matrices $\bm{X}$ and E can be augmented with arbitrary node and edge features deemed useful (indicated with the $+$ superscript in Figure 2(a)).
+In addition to the two matrices, a further vector $\bm{u}$ is used as input, containing the current timestep $t$ along with optional graph-level features (such as the overall molecular weight).
+All these data are augmented with the guidance $\bm{y}$ and processed by the stack of graph transformer layers.
 The graph transformer is displayed in Figure 2(b): at a high level, it resembles a standard transformer architecture, with a self attention module followed by dropout, residual connection, layer norm, and a feedforward stack.
 The self attention layer (shown in Figure 2(c)) first processes the (augmented) $\bm{X}$ matrix through the standard outer product and scaling typical of attention layers.
-Then, the result is used in conjunction with the edge tensor E as input to a FiLM layer (Perez et al.
-2018) to produce unnormalized attention scores.
+Then, the result is used in conjunction with the edge tensor E as input to a FiLM layer [15] to produce unnormalized attention scores.
 The scores are passed through a softmax layer, combined once again with the values $\bm{X}$.
@@
 The vector $\bm{u}$ is instead processed differently, as shown visually in the rightmost part of Figure 2(c).
-Specifically, $\bm{X}$ and E are first passed through independent PNA layers (Corso et al.
-2020) and then summed together with a linear projection of $\bm{u}$ to form the final representation.
-Loss function.
+Specifically, $\bm{X}$ and E are first passed through independent PNA layers [3] and then summed together with a linear projection of $\bm{u}$ to form the final representation.
+3.0.4 Loss function.
 During training, FreeGress optimizes the following loss function for a given time step $t$:
-$\delta\sum_{1\leq i\leq n}\mathrm{NLL}\left(\bm{x}_{i}^{0},f_{\theta}(\bm{x}_{i}^{t})\right)+\gamma\sum_{1\leq i,j\leq n}\mathrm{NLL}\left(\bm{e}_{ij}^{0},f_{\theta}(\bm{e}_{ij}^{t})\right),$
-which is the negative log-likelihood (NLL) between the actual node (resp.
+$\sum_{1\leq i\leq n}\mathrm{CE}\left(\bm{x}_{i}^{0},f_{\theta}(\bm{x}_{i}^{t})\right)+\gamma\sum_{1\leq i,j\leq n}\mathrm{CE}\left(\bm{e}_{ij}^{0},f_{\theta}(\bm{e}_{ij}^{t})\right),$ (12)
+which is the cross-entropy (CE) between the actual node (resp.
 edge) types and those predicted by the neural network described above.
-Parameters $\delta$ and $\gamma$ adjust the preference towards nodes or edges prediction.
-Conditioned Inference
-At each reverse process iteration, FreeGress generates a slightly denoised graph by first computing from the network $\log{p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{y})}$, the conditioned log probability, and $\log{p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{\bar{y}})}$, the unconditioned one.
-The final probability is obtained by exponentiating the sum of the two terms while giving an arbitrary (usually larger) weight $s$ to the conditioned probability:
-$\log{p(\bm{x}^{0}_{i}|G^{t})}+(s+1)\,(\log{p(\bm{x}^{0}_{i}|G^{t},\bm{y})}-\log{p(\bm{x}^{0}_{i}|G^{t})}),$
-where we omit exponentiation for conciseness, which corresponds to $f_{\theta}(\bm{x}^{0}=\bm{x}|G^{t},\bm{y})$ in Eq.
-12.
-Then, the posteriors $q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})$ are computed $\forall\bm{x}\in\mathcal{X}$ and summed as in Eq.
-9.
+$\gamma$ is a hyper-parameter which adjusts the preference towards edges prediction.
+When Equation 11 is employed, CE is replaced by the negative log-likelihood loss.
+3.0.5 Conditioned Inference.
+At each reverse process iteration, FreeGress generates a slightly denoised graph by first computing from the network $p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{y})$, the conditioned probability, and $p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{\bar{y}})$, the unconditioned one.
+The final probability is obtained by summing the two terms while giving an arbitrary (usually larger) weight $s$ to the conditioned probability, using Equation 10.
+Then, the posteriors $q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})$ are computed $\forall\bm{x}\in\mathcal{X}$ and summed as in Equation 7.
 This procedure is executed for each atom in the graph, and the resulting probabilities are sampled to obtain the nodes of the graph $G^{t-1}$.
 The process is symmetric for the graph edges.
-3.2 Graph size inference
-One limitation of diffusion models as a whole is the fact that the size (for graphs specifically, the number of nodes) of a sample cannot change during the denoising process.
-Some studies consider the possibility of inserting and removing elements from a generated, mono-dimensional sequence (Johnson et al.
-2021) but, to the best of our knowledge, there are no similar works for graphs.
+3.1 Conditioning on the number of nodes
+A general limitation of diffusion models is the fact that the size of a sample (for graphs specifically, the number of nodes) cannot change during the denoising process.
+Some studies consider the possibility of inserting and removing elements from a generated, mono-dimensional sequence [12] but, to the best of our knowledge, there are no similar works for graphs.
 In the original implementation of DiGress, the number of nodes the generated graph will have is sampled from the marginal distribution computed from the training and validation sets, respectively.
@@
 4 Experiments
-Here, we detail the experimental analysis to evaluate FreeGress on property targeting tasks, that is, generating compounds which meet pre-specified molecular properties.
-4.1 Dataset
+Here, we detail the experimental analysis to evaluate FreeGress on property targeting tasks by generating compounds that meet pre-specified properties.
+4.1 Datasets
 Our experiments were conducted on QM9, a dataset of 133k small molecules (up to 9 heavy atoms) which was also part of the evaluation of DiGress, and ZINC-250k, a collection of 250k drug-like molecules selected from the ZINC dataset.
-For the latter, the molecules were first preprocessed by removing non-neutral charges and stereochemistry information, reducing the dataset size to 204k molecules.
+For the latter, the molecules were first preprocessed by removing stereochemistry information.
+We have also removed scarcely present non-neutrally-charged atoms, leaving only N+ and O-.
+The latter have been treated as standalone atom types instead of their neutrally-charged counterparts.
+The final dataset size was consequently reduced to 228k molecules.
+Further statistics about the datasets are provided in the Supplementary Section 1.
 4.2 Metrics and targets
-All the experiments were performed using the following scheme.
+All the experiments were performed using the following setup.
 After we trained each model, we randomly sampled 100 molecules from the dataset and computed the desired target properties as vectors $\bm{y}$, which were used as conditioning vectors.
 Then, we used the vectors 10 times each to perform conditioned generation, for a total of 1000 generated molecules for each model.
-We computed the desired properties on the generated samples [^1] Using packages such as RDKit (Landrum 2016) and psi4 (Turney et al.
-2012) and compared the results with the properties of the original samples.
+We computed the desired properties on the generated samples [^1] Using packages such as RDKit [13] and psi4 [23] and compared the results with the properties of the original samples.
 The metric chosen for the comparison is the Mean Absolute Error (MAE) between the target properties and the properties of the generated molecules:
@@
 where $\bm{y}_{i}$ are the target properties of the $i$-th molecule from the dataset and $\widehat{\bm{y}}_{i,j}$ is the target property of the $j$-th molecule generated using the properties of the $i$-th molecule as guide.
-The targeted properties were dipole moment $\mu$ and the Highest Occupied Molecular Orbital (HOMO) for QM9; for ZINC-250k, we targeted log-partition coefficient (LogP) and Molecular Weight (MW).
-Model Hyper-parameters Extra features HOMO $\downarrow$ $\bm{\mu}$downarrow HOMO + $\bm{\mu}$downarrow
-DiGress Not provided No $0.56\pm 0.01$ $0.81\pm 0.04$ $0.87\pm 0.03$
-FreeGress $p_{uncond}=0,s=5$ No $0.50\pm 0.05$ $0.99\pm 0.08$ $0.74\pm 0.06$
-$p_{uncond}=0,s=5$ Yes $0.34\pm 0.03$ $0.95\pm 0.07$ $0.62\pm 0.04$
-$p_{uncond}=0.1,s=5$ No $0.30\pm 0.04$ $0.60\pm 0.06$ $0.59\pm 0.07$
-$p_{uncond}=0.1,s=5$ Yes $\mathbf{0.28\pm 0.04}$ $\mathbf{0.54\pm 0.06}$ $\mathbf{0.53\pm 0.07}$
-$p_{uncond}=0.2,s=5$ No $0.31\pm 0.05$ $0.64\pm 0.07$ $0.62\pm 0.06$
-$p_{uncond}=0.2,s=5$ Yes $0.36\pm 0.10$ $0.57\pm 0.07$ $0.55\pm 0.06$
-Model Hyper-parameters LogP MW
-MAE $\downarrow$ Valid $\uparrow$ MAE $\downarrow$ Valid $\uparrow$
-DiGress $\lambda=50$ $1.15\pm 0.09$ 75% $73.40\pm 7.10$ 75%
-$\lambda=500$ $0.60\pm 0.05$ 68% $60.30\pm 6.10$ 73%
-$\lambda=1000$ $0.49\pm 0.05$ 65% $61.10\pm 5.90$ 69%
-$\lambda=3000$ $0.52\pm 0.13$ 39% $64.10\pm 5.90$ 59%
-FreeGress $p_{uncond}=0.1,s=2$ $\mathbf{0.15\pm 0.02}$ 85% $15.73\pm 2.74$ 58%
-$p_{uncond}=0.1,s=2$, NI $0.16\pm 0.01$ 84% $\mathbf{15.18\pm 2.71}$ 75%
-Table 1: Results on the QM9 (top) and ZINC-250k (bottom) datasets.
-Property targeting has been evaluated using Mean Absolute Error (see Section 4.2).
-HOMO: Highest Occupied Molecular Orbital; $\mathbf{\mu}$: dipole moment; $\lambda$: weight of the classier gradient for CB guidance; $s$: weight of conditioned inference for FreeGress; $p_{uncond}$: conditional dropout parameter for FreeGress; Valid: percentage of valid molecules among those sampled; NI: node inference (see Section 3.2).
+The targeted properties were the dipole moment $\mu$ and the Highest Occupied Molecular Orbital (HOMO) for QM9; for ZINC-250k, we targeted the Log-Partition coefficient (LogP), and the Quantitative Estimation of Drug-likeness (QED).
+The proposed node inference method was evaluated on the ZINC-250k dataset targeting Molecular Weight (MW).
 4.3 Experimental details
-QM9.
-On QM9, we trained a set of FreeGress variants against the published results of DiGress.
-In the original article, DiGress was trained for the property targeting task without additional features, and hyper-parameters such as $\lambda$ (which regulates the impact of the regressor’s gradient) were not provided.
-FreeGress instances were trained with $s=5$, and with $p_{uncond}=\{0,0.1,0.2\}$ to study the effect of conditional dropout on the generative process.
-Moreover, we also trained different variants with and without additional features for $\bm{X}$ and $\bm{u}$, to understand if they provided any positive impact.
-Additional graph-wise features considered for this study were number of $k$-cycles with $k\in\{1,\ldots,6\}$, number of connected components, first 5 non-zero eigenvalues, molecular weight; among node-wise features, we considered the atom’s valency, the number of $k$-cycles a node belongs to, and an estimation of the biggest connected component size.
-FreeGress used similar hyper-parameters (5 graph transformer layers; embedding size of 256 for $\bm{X}$ and 128 for both E and $\bm{u}$; 8 attention heads; $T=500$, 1200 training epochs with batch size 512; Amsgrad optimizer (Reddi, Kale, and Kumar 2018) with learning rate of $2^{-4}$ and weight decay of $1^{-12}$; $\gamma=2\delta$) as DiGress for a fair comparison.
+4.3.1 QM9 for $\mu$/HOMO targeting.
+We trained and compared a set of FreeGress variants against a set of DiGress models, including an unconditional variant as baseline.
+To the best of our knowledge, DiGress is the only model that allows for property targeting at the time of writing.
+FreeGress instances were trained with $s=\{1,3,5\}$, and with $\rho=\{0,0.1,0.2\}$ to study the effect of conditional dropout on the generative process.
+We have also experimented variants with additional features and $\rho=\{0,0.1\}$; in this case, $\rho=0.2$ was excluded as it did not improve on the former in preliminary trials.
+For DiGress, we evaluated the performance using $\lambda=\{200,400,600,800\}$.
+To ensure a fair comparison, we trained both DiGress and FreeGress with matching architectural design: specifically, we used 5 graph transformer layers; embedding size of 256 for $\bm{X}$ and 128 for both E and $\bm{u}$; 8 attention heads; $T=500$, 1200 training epochs with batch size 512; Amsgrad optimizer [18] with learning rate of $2^{-4}$ and weight decay of $1^{-12}$; $\gamma=2$.
 Importantly, since FreeGress does not require an auxiliary regressor, all the variants are trained with half the number of parameters than DiGress (4.6 million vs.
 9.2 million, approximately).
-All models required approximately seven hours to train on a nVidia A100 with 80GB of VRAM.
-ZINC-250k.
-For this dataset, we used a slightly different experimental setup since DiGress was not evaluated on this dataset previously.
-Specifically, we trained from scratch both DiGress and FreeGress with 12 layers and 1000 training epochs with batch size 256, while keeping the other hyper-parameters the same as for QM9.
+A single training required up to 12 hours on an nVidia V100 with 16GB of VRAM.
+4.3.2 ZINC-250k for logP/QED targeting.
+For these experiments we used a slightly different setup.
+Specifically, we trained from scratch both DiGress and FreeGress with 12 layers, 1000 training epochs, batch size 256, and doubling the embedding size of $\bm{y}$, while keeping the other hyper-parameters the same.
 We used larger models since ZINC-250k includes bigger (in terms of number of atoms) and more diverse molecules.
-Training required approximately 100 hours for 1000 epochs each on the same machine as above.
-While FreeGress variants used approximately 16 million parameters, DiGress used approximately 32 million including the property regressor.
-We trained without additional features, since preliminary experiments have revealed that they had little to no influence on the performance, while at the same time largely increasing the training time.
-For DiGress, we evaluated different variants alternatives with $\lambda=\{50,500,1000,3000\}$.
-For FreeGress, we fixed $s=2$ and $p_{uncond}=0.1$ and evaluated one variant with the inference method proposed in Section 3.2, and one without.
-In the former case, $p_{\varepsilon}(n|\bm{y})$ was parameterized as a neural network with 2 layers with 512 units each, and ReLU activations.
-Input $\mu$: 2.9326Estimated $\mu$: 2.8836
-Input $\mu$: 4.1600Estimated $\mu$: 4.0071
-Input HOMO: -6.7892Estimated HOMO: -6.6912
-Input $\mu$: 2.9326Estimated $\mu$: 2.8836
-Input $\mu$: 4.1600Estimated $\mu$: 4.0071
-Input HOMO: -6.7892Estimated HOMO: -6.6912
-Input HOMO: -7.4478Estimated HOMO: -7.3225
-Figure 2: Curated molecules from the QM9 dataset.
+In this case, we did not perform experiments with additional features since they increased training time by a prohibitive amount.
+FreeGress variants required approximately 16 million parameters, while DiGress used approximately 32 million including the property regressor.
+Training the various models required approximately 5 days each on an nVidia A100 with 80GB of VRAM.
+4.3.3 ZINC-250k for MW targeting.
+These experiments evaluate the node inference method proposed in Section 3.1.
+As such, we targeted MW since it trivially depends on the number of nodes.
+The setup is similar to the logP/QED experiments, with some slight differences.
+In this case only, we have managed to train FreeGress with additional features and $\rho\in\{0,0.1\}$ ($\rho=0.2$ was excluded since it did not bring significant improvements in early trials).
+The node inference model $p_{\varepsilon}(n|\bm{y})$ was parameterized as a neural network with 2 layers with 512 units each and ReLU activations.
+Table 1: Results on the QM9 dataset. First two columns are single conditioning, last column is multiple conditioning on $\mu$ and HOMO.
+$\mu$ HOMO $\mu$+HOMO
+MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$
+Unconditional $1.68\pm 0.15$ $91.5\%$ $0.95\pm 0.10$ $91.5\%$ $1.32\pm 0.1$ $91.5\%$
+DiGress
+$\lambda=200$ $1.06\pm 0.13$ $87.0\%$ $0.70\pm 0.07$ $91.7\%$ $1.13\pm 0.08$ $88.7\%$
+$\lambda=400$ $0.95\pm 0.07$ $86.5\%$ $0.65\pm 0.07$ $91.7\%$ $1.08\pm 0.09$ $87.2\%$
+$\lambda=600$ $0.80\pm 0.07$ $82.5\%$ $0.61\pm 0.07$ $91.2\%$ $1.11\pm 0.10$ $86.7\%$
+$\lambda=800$ $0.86\pm 0.08$ $76.7\%$ $0.52\pm 0.05$ $90.5\%$ $1.12\pm 0.11$ $87.1\%$
+FreeGress
+$\rho=0$ $1.03\pm 0.10$ $86.7\%$ $0.45\pm 0.05$ $92.0\%$ $0.83\pm 0.11$ $82.3\%$
+$\rho=0.1,s=1$ $1.06\pm 0.12$ $88.3\%$ $0.51\pm 0.07$ $92.3\%$ $0.87\pm 0.09$ $86.2\%$
+$\rho=0.1,s=3$ $0.74\pm 0.08$ $83.7\%$ $0.32\pm 0.04$ $90.1\%$ $0.68\pm 0.09$ $83.4\%$
+$\rho=0.1,s=5$ $0.67\pm 0.10$ $77.7\%$ $0.40\pm 0.08$ $92.9\%$ $0.77\pm 0.13$ $77.0\%$
+$\rho=0.2,s=1$ $1.03\pm 0.10$ $89.1\%$ $0.47\pm 0.05$ $91.6\%$ $0.84\pm 0.10$ $84.3\%$
+$\rho=0.2,s=3$ $0.83\pm 0.10$ $82.1\%$ $0.32\pm 0.04$ $93.8\%$ $0.80\pm 0.18$ $72.0\%$
+$\rho=0.2,s=5$ $0.74\pm 0.11$ $75.5\%$ $0.44\pm 0.07$ $93.3\%$ $0.87\pm 0.19$ $72.7\%$
+FreeGress+extra features
+$\rho=0$ $0.90\pm 0.10$ $90.1\%$ $0.36\pm 0.04$ $94.2\%$ $0.65\pm 0.07$ $93.1\%$
+$\rho=0.1,s=1$ $0.90\pm 0.10$ $90.0\%$ $0.35\pm 0.03$ $92.6\%$ $0.65\pm 0.07$ $92.5\%$
+$\rho=0.1,s=3$ $0.60\pm 0.07$ $85.2\%$ $0.27\pm 0.04$ $92.1\%$ $0.54\pm 0.07$ $90.0\%$
+$\rho=0.1,s=5$ $0.52\pm 0.07$ $81.5\%$ $0.40\pm 0.08$ $91.1\%$ $0.59\pm 0.11$ $84.9\%$
 5 Results
-The experimental results are presented in Table 1.
-Novelty and uniqueness scores are not reported as they approached 100% in almost all cases.
-On the QM9 dataset (top table), we observe that FreeGress outperforms DiGress in almost every variant tested.
-Interestingly enough, even variants of FreeGress with $p_{uncond}=0$, i.e.
-those that use the guide without conditional dropout, achieve comparable performance with DiGress, with the only exception of the dipole moment $\mu$.
-Overall, the best performing variant of FreeGress ($p_{uncond}=0.1$, $s=5$, additional features) improves with respect of DiGress by 50% on the HOMO task, by 33% on the $\mu$ task, and by 39% on the combination of both.
-Similar improvements are achieved with slightly different combinations of the hyper-parameters, showing that CF guidance is robustly a better alternative to CB guidance for conditioned generation.
-Also, one should take into account that CF guidance does not require to train a separate auxiliary regressor, which saves up computational and memory resources.
-Four conditioned samples (two for each property targeting task) generated from the QM9 dataset are shown in Figure 2.
-The analysis of performance in the ZINC-250k provides some additional interesting insights.
-Firstly, we observe a drop in validity of DiGress whenever too much weight is put on the auxiliary regressor by increasing $\lambda$.
-This is a clear signal that the gradient of the regressor is not providing useful directions in chemical space, which validates our belief that CB guidance, as formulated in DiGress, is based on unrealistic assumptions for the chemical domain.
-Secondly, we observe that FreeGress outperforms again DiGress on the two tasks by a great margin.
-Specifically, on the LogP task, FreeGress achieves a 69% improvement in MAE with respect to the best DiGress variant, while at the same time improving validity by 31%.
-On the MW task, the improvement in MAE is of about 79% with respect to the best DiGress variant.
-Thirdly, we observe that node inference has not a decisive impact in the LogP task.
-However, in the MW task, it improves the validity rate from 58% to 75% (on par with the best DiGress variant), while also slightly improving in terms of MAE.
-This aligns with our intuition: in fact, molecules with higher molecular weight are in general more likely to have more atoms than molecules with lower molecular weight.
-Thus, it makes sense, in this particular task, to having a mechanism to infer the number of nodes from the guide to bias the sampling process.
-Figure 2 shows 4 exemplay molecules (2 for each task) generated from the ZINC-250k dataset.
+5.0.1 QM9.
+The experimental results on QM9 are presented in Table 1, where we report the MAE and the chemical validity of the generated molecules.
+Uniqueness scores are not reported as they approached 100% in almost all cases, while the novelty is not a relevant metric on QM9 as the dataset is an enumeration of all the valid molecules satisfying a particular set of constraints.
+We observe that FreeGress variants, and especially those that leverage additional features, outperform DiGress variants in the vast majority of cases.
+Interestingly enough, even FreeGress variants with $\rho=0$, i.e.
+those that use the guide without conditional dropout, are competitive with DiGress.
+Overall, FreeGress convincingly outperforms DiGress even when the additional features are not used (by 16%, 38% respectively on the $\mu$ and HOMO properties, and by 37% when conditioning on both).
+When targeting $\mu$+HOMO in conjunction, the gap in performance indicates that FreeGress is better suited for multiple conditioning.
+As we show in Supplementary Table 1, this might be attributed to the property regressor of DiGress, which struggles to generate molecules conditioned on both properties, while FreeGress does not suffer from this issue.
+Overall, we observe a slight inverse correlation between MAE and validity, with one tending to decrease as the other improves.
+Table 2: Results on the ZINC-250k dataset. First two columns are single conditioning, last column is multiple conditioning on LogP and QED.
+LogP QED LogP+QED
+MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$
+Unconditional $1.52\pm 0.12$ $86.1\%$ $0.15\pm 0.01$ $86.1\%$ $0.83\pm 0.06$ $86.1\%$
+DiGress
+$\lambda=200$ $0.83\pm 0.08$ $77.6\%$ $0.15\pm 0.01$ $85.8\%$ $0.49\pm 0.03$ $83.8\%$
+$\lambda=400$ $0.74\pm 0.08$ $74.6\%$ $0.15\pm 0.01$ $85.1\%$ $0.40\pm 0.02$ $78.2\%$
+$\lambda=600$ $0.76\pm 0.10$ $69.6\%$ $0.14\pm 0.01$ $84.5\%$ $0.38\pm 0.02$ $79.3\%$
+$\lambda=800$ $0.92\pm 0.12$ $65.1\%$ $0.14\pm 0.01$ $83.8\%$ $0.35\pm 0.01$ $75.5\%$
+FreeGress
+$\rho=0$ $0.22\pm 0.02$ $80.9\%$ $0.06\pm 0.01$ $84.7\%$ $0.15\pm 0.01$ $82.3\%$
+$\rho=0.1,s=1$ $0.22\pm 0.02$ $84.7\%$ $0.06\pm 0.01$ $83.8\%$ $0.16\pm 0.01$ $83.1\%$
+$\rho=0.1,s=2$ $0.17\pm 0.01$ $84.9\%$ $0.04\pm 0.01$ $84.9\%$ $0.12\pm 0.01$ $80.7\%$
+$\rho=0.1,s=3$ $0.18\pm 0.02$ $82.2\%$ $0.05\pm 0.01$ $85.6\%$ $0.12\pm 0.01$ $77.5\%$
+$\rho=0.2,s=1$ $0.22\pm 0.02$ $87.0\%$ $0.07\pm 0.01$ $84.2\%$ $0.17\pm 0.02$ $78.6\%$
+$\rho=0.2,s=2$ $0.19\pm 0.02$ $83.2\%$ $0.05\pm 0.01$ $82.8\%$ $0.13\pm 0.01$ $76.8\%$
+$\rho=0.2,s=3$ $0.16\pm 0.01$ $81.2\%$ $0.05\pm 0.01$ $80.8\%$ $0.13\pm 0.02$ $73.1\%$
+5.0.2 ZINC-250k
+The experimental results of targeting logP and QED (alone and in conjunction) are presented in Table 2.
+Uniqueness and novelty scores are not reported as they approached 100% in almost all cases.
+Similarly to QM9, we observe that FreeGress outperforms DiGress on all tasks by a wide margin.
+For example, on the LogP task, FreeGress with $\rho=0.2,s=3$ achieves a 78% decrease in MAE with respect to the best DiGress variant ($\lambda=400$), while at the same time improving validity.
+On the QED task, FreeGress with $\rho=0.1,s=2$ achieves a 71% MAE decrease with respect to the best DiGress variant ($\lambda=600$), while keeping the same validity rates.
+When conditioning on LogP and QED in conjunction, FreeGress with $\rho=0.1,s=2$ decreases MAE by 65% in MAE compared to the best DiGress variant ($\lambda=800$) while slightly improving validity by 5%.
+Overall, the results empirically confirm that FreeGress performs better than DiGress regardless of the dataset where it is applied and the properties that is tasked to target.
+Table 3: Results of applying the proposed node inference method $p_{\varepsilon}(n\,|\,\bm{y})$ to FreeGress while targeting Molecular Weight on the ZINC-250k dataset.
+The last column reports the improvement against a “no inference” FreeGress variant.
+No inference $p_{\varepsilon}(n\,|\,\bm{y})$ Improvement
+MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$
+DiGress
+[…]
+$\lambda=400$ $64.75$ p m 6.05 79.3% $22.51$ p m 2.62 38.4% -65.2% -51.6%
+$\lambda=600$ $64.67$ p m 6.17 80.2% $21.70$ p m 3.00 41.2% -66.4% -48.6%
+$\lambda=800$ $62.96$ p m 6.21 77.8% $20.92$ p m 2.90 40.4% -66.8% -48.1%
+FreeGress
+$\rho=0$ $\bm{7.08\pm 1.19}$ 54.5% $7.25$ p m 1.81 75.7% +2.4% +38.9%
+$\rho=0.1,s=1$ $14.93$ p m 2.27 60.9% $8.22$ p m 0.77 81.4% -44.9% +33.7%
+$\rho=0.1,s=2$ $11.37$ p m 2.81 50.8% $8.96$ p m 1.93 79.7% -21.2% +56.9%
+$\rho=0.1,s=3$ $9.33$ p m 1.87 46.5% $8.03$ p m 1.02 66.6% -13.9% +43.2%
+$\rho=0.2,s=1$ $20.00$ p m 5.47 61.3% $13.22$ p m 1.51 81.3% -33.9% +32.6%
+$\rho=0.2,s=2$ $13.69$ p m 3.19 53.9% $13.37$ p m 1.47 78.4% -2.3% +45.5%
+$\rho=0.2,s=3$ $15.61$ p m 4.08 48.4% $14.28$ p m 2.12 75.0% -8.5% +54.9%
+FreeGress+extra features
+$\rho=0$ $12.66$ p m 4.99 60.1% $5.13$ p m 0.38 $\bm{86.1\%}$91.85% +43.25%
+$\rho=0.1,s=1$ $43.27$ p m 9.35 82.1% $3.52$ p m 0.34 83.7% -91.87% +1.95%
+$\rho=0.1,s=2$ $48.25$ p m 11.03 73.3% $\bm{3.34\pm 0.31}$ 84.7% -93.08% +15.55%
+$\rho=0.1,s=3$ $47.67$ p m 13.26 63.2% $3.35$ p m 0.40 81.2% -92.97% +28.48%
+5.0.3 Effect of node inference in ZINC-250k
+Table 3 summarizes the results on applying node inference on the MW targeting task.
+The “no inference” scenario (first column) indicates that FreeGress obtains lower MAEs than DiGress, regardless of whether additional features are added or not.
+However, this happens at the expense of validity, which is lower for FreeGress than DiGress.
+Adding node inference (central column) allows FreeGress to generate valid molecules at the same rate than DiGress, but with better MAE.
+Moreover, the addition of node inference is beneficial to FreeGress itself, as quantified by the third column which measures row-wise improvement.
+Surprisingly, we observe that applying node inference to DiGress decreases MAE by approximately 71% across experiments, but at the same time completely disrupts chemical validity.
+In contrast, adding node inference to FreeGress causes a good decrease in MAE and a steep increase in validity (up to 57%).
+Overall, the best improvement is given by the combination of node inference and additional molecular features.
+Input $\mu$: 0.0603
+Est. $\mu$: 0.0463
+Input $\mu$: 4.2338
+Est. $\mu$: 4.1238
+Input HOMO: -6.8083
+Est. HOMO: -6.8604
+Input HOMO: -7.4559
+Est. HOMO: -7.4632
+Input LogP: -0.9834
+Est. LogP: -0.9246
+Input QED: 0.5034
+Est. QED: 0.5297
+Input MW: 159.05
+Est. MW: 155.03
+Input MW: 401.13
+Est. MW: 405.10
+Figure 3: Curated molecules from the QM9 (top row) and ZINC-250k (bottom row). The input conditioning value and the one estimated (Est.) after generation are displayed below each molecule.
+5.0.4 Curated samples.
+In Figure 3 we show some curated samples from the targeting properties under study under study to add visual evidence that the model is operating correctly and it is able to target different properties with success.
+Additional random samples are shown in the Supplementary material.
 6 Conclusions
-This work has developed and implemented FreeGress, a discrete DDPM for graphs with CF guidance, which allows to generate molecules complying to a pre-specified set of desired chemical properties.
-Through experiments, we have shown that CF guidance allows to generate better (more tailored to the specification) molecules than CB guidance without sacrificing chemical validity.
-Of course, there are still limitations to overcome, which we intend to address in future studies: for example, even if CF guidance allows to achieve good validity rates, the model is still far from achieving perfect validity.
-This might be addressed by designing the forward and reverse processes to use conditional transition matrices.
-Another way to increase the percentage of valid molecules generated without overfitting may be to train a model to work with molecular fragments instead of atoms, as it would likely reduce the amount of points in the molecular graph where an illegal structure is generated.
-We also have shown that inferring the number of nodes from the guide might be of help to generate graphs in some settings, especially when there is reason to believe that the guide is correlated with the graph’s size.
-As future work, we intend to validate this finding also in the more general setting of arbitrary graph generation.
-Input LogP: 0.8472Estimated LogP: 0.8590
-Input LogP: 2.1313Estimated LogP: 2.0127
-Input MW: 401.13Estimated MW: 411.06
-Input LogP: 0.8472Estimated LogP: 0.8590
-Input LogP: 2.1313Estimated LogP: 2.0127
-Input MW: 401.13Estimated MW: 411.06
-Input MW: 195.02Estimated MW: 196.18
-Figure 3: Curated molecules from the ZINC-250k dataset.
+We introduced FreeGress, a discrete DDPM for graphs with CF guidance, which allows to generate molecules complying with a pre-specified set of desired chemical properties.
+Through extensive experiments, we have shown that CF guidance allows to generate better (more tailored to the specification) molecules than CB guidance without sacrificing chemical validity.
+Additionally, we have implemented a form of learned node inference, and shown that inferring the number of nodes from the guide helps the generation whenever the molecular size is connected to the target property.
+In future studies, we intend to tackle some of the current limitations of FreeGress to expand its applicability.
+A first research direction is to improve chemical validity designing the forward and reverse processes to use conditional transition matrices, thus constraining the intermediate noising and denoising steps to happen in valid chemical space.
+Another promising direction is to work with molecular fragments rather than atoms, as it would likely reduce the opportunities to generate an invalid intermediate structure.
+We intend to validate these findings also in the more general setting of arbitrary graph generation.
+{credits}
+6.0.1 Acknowledgements
+Research partly funded by PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - "FAIR - Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI", funded by the European Commission under the NextGeneration EU programme.
+[…]
+The authors have no competing interests to declare that are relevant to the content of this article.